# Vision-Language Models Represent Darker-Skinned Black Individuals as More Homogeneous than Lighter-Skinned Black Individuals

### v1

@@ Abstract
Vision-Language Models (VLMs) combine Large Language Model (LLM) capabilities with image processing, enabling tasks like image captioning and text-to-image generation. Yet concerns persist about their potential to amplify human-like biases, including skin tone bias. Skin tone bias, where darker-skinned individuals face more negative stereotyping than lighter-skinned individuals, is well-documented in the social sciences but remains under-explored in Artificial Intelligence (AI), particularly in VLMs. While well-documented in the social sciences, this bias remains under-explored in AI, particularly in VLMs. Using the GAN Face Database, we sampled computer-generated images of Black American men and women, controlling for skin tone variations while keeping other features constant. We then asked VLMs to write stories about these faces and compared the homogeneity of the generated stories. Stories generated by VLMs about darker-skinned Black individuals were more homogeneous than those about lighter-skinned individuals in three of four models, and Black women were consistently represented more homogeneously than Black men across all models. Interaction effects revealed a greater impact of skin tone on women in two VLMs, while the other two showed nonsignificant results, reflecting known stereotyping patterns. These findings underscore the propagation of biases from single-modality AI systems to multimodal models and highlight the need for further research to address intersectional biases in AI.
@@ 1 Introduction
1.1 Homogeneity Bias in AI Models
Another form of stereotyping documented in AI models is homogeneity bias—the tendency of AI systems to represent certain groups as more uniform than others. Rooted in social psychology, this phenomenon relates to the out-group homogeneity effect, where individuals perceive their own group (in-group) as more diverse than other groups (out-group) (Linville et al., 1989; Quattrone and Jones, 1980).
Lee et al. (2024) explored whether a state-of-the-art LLM reproduces similar biases, examining how group status affects the homogeneity of group representations. They found that ChatGPT generated more uniform texts for racial/ethnic minority groups and women compared to the racial/ethnic majority group in the US (i.e., White Americans) and men. They suggested that disproportionate representation and stereotypical portrayals of marginalized groups in training data may lead LLMs to replicate this pattern of human stereotyping. Similarly, Cheng et al. (2023) found that texts about marginalized groups often exaggerate defining characteristics, resulting in more caricature-like representations.
@@ 1.1 Homogeneity Bias in AI Models
1.2 Skin Tone Bias
Skin tone bias refers to prejudiced attitudes or discriminatory behaviors directed toward individuals within a racial group based on their skin tone (Maddox, 2004; Maddox and Gray, 2002). Research in this area has predominantly focused on perceptions of Black individuals, showing that darker-skinned Black individuals are more likely to be associated with negative traits compared to lighter-skinned Black individuals (e.g., Berumen Martinez, 2021; Stepanova and Strube, 2018; Kahn and Davies, 2011; Lynn, 2002; Maddox and Gray, 2002).
This bias also has significant real-world consequences. For example, darker-skinned Black individuals are receive lower ratings and fewer job offers in hiring scenarios (Wade et al., 2004; Harrison and Thomas, 2009), achieve lower levels of educational attainment and income (Keith and Herring, 1991), experience greater racial discrimination (Klonoff and Landrine, 2000), and report higher levels of mental distress due to discrimination (Gleiberman et al., 1995). Overall, skin tone plays a critical role in shaping both perceptions and life outcomes for Black individuals in the United States.
@@ 1.3 This Work
While substantial research in the social sciences has examined the impact of skin tone on social perceptions and real-world outcomes, relatively few studies have addressed skin tone bias in Artificial Intelligence (AI) models, particularly in Vision-Language Models (VLMs). Seminal work by Buolamwini and Gebru (2018) revealed that commercial gender-classification systems perform significantly better for lighter-skinned individuals. While much subsequent research has investigated the impact of skin tone on machine learning model performance (e.g., Groh et al., 2024; Kinyanjui et al., 2019), the topic of skin tone bias in newer generative models (e.g., VLMs) remains relatively under-explored.
Figure 1: Summary of the experimental setup. We collect 50-word stories about Black individuals differing in skin tone using four state-of-the-art Vision-Language Models. After encoding these stories into sentence embeddings, we compare the pairwise similarity of the embeddings using mixed-effects models.
Skin tone bias in Vision-Language Models (VLMs) carries significant implications. As VLMs increasingly contribute to the production of social information and creative content, they may generate more stereotypical texts in response to images of darker-skinned individuals compared to lighter-skinned individuals. Research shows that repeated exposure to stereotypes in social information can reinforce societal biases (Park et al., 2007; Saleem et al., 2017), suggesting that if VLMs perpetuate skin tone bias, they could exacerbate the very inequities that led to the bias in their training data. Moreover, as VLMs and other generative models are increasingly deployed in decision-making contexts, such biases could amplify discrimination in real-world outcomes.
We explored the topic of skin tone bias in VLMs through the lens of homogeneity bias, investigating how skin tone is associated with homogeneity of group representations. Our study was informed by hypotheses and research questions grounded in the social psychology literature on skin tone bias. Based on extensive evidence of the role of skin tone in stereotyping, we hypothesized that stories about darker-skinned Black individuals would exhibit greater homogeneity than those about lighter-skinned Black individuals. To test this, we encoded VLM-generated stories into sentence embeddings and measured their pairwise cosine similarity, following the method introduced by Lee et al. (2024). A mixed-effects model was then fitted with skin tone, derived from the images, as the only predictor. Additionally, we investigated the interaction between skin tone and gender as a research question. Prior studies suggest that skin tone effects may differ across genders, with stronger influences on attractiveness perceptions among women (Hill, 2002) and more pronounced performance disparities found in commercial gender-classification systems within women compared to men (Buolamwini and Gebru, 2018).
@@ 2.1 Image Stimuli
We sampled ten images of Black American men and women from the publicly available GAN Face Database (GANFD) [^1] https://osf.io/7auyw, which features realistic, computer-generated faces. The database includes sets of images representing the same fictional individuals, with manipulations applied to vary facial features associated with perceived race. To ensure consistency, we used cropped images with a uniform grey background that contained only the face.
Figure 2: Two face stimulus pairs representing (A) Black men and (B) Black women. In each pair, the left image depicts a lighter-skinned individual, while the right shows a darker-skinned individual, generated from the same set.
From the GANFD, we focused on facial stimuli that were primarily identified as Black or multiracial. [^2] Multiracial faces were included due to an insufficient number of faces rated as Black. Within each set, faces with higher perceived Blackness ratings, measured by the question, “How much does this person look Black or African American? (0-100),” were assigned to the darker-skinned condition, while those with lower perceived Blackness scores were assigned to the lighter-skinned condition. As shown in Figure 2, faces with higher perceived Blackness typically had darker skin tones, while other facial features remained largely consistent across the skin tone conditions. This approach allowed us to manipulate skin tone while controlling for other features associated with perceived Blackness, such as nose width or eyebrow thickness (Ma et al., 2018).
@@ 2.2 Selection of Vision-Language Models and Writing Prompts
We collected data using a collection of Vision-Language Models including two proprietary VLMs—GPT-4o mini, GPT-4 Turbo—and two open-source VLMs—Llama-3.2-11B-Vision-Instruct (“Llama-3.2") (Grattafiori et al., 2024) and BLIP-3 (xgen-mm-phi3-mini-instruct-r-v1) (Xue et al., 2024). We accessed the proprietary models using the OpenAI API and the open-source models by downloading the model weights and running inferences on them locally.
The models were given the following writing prompt, “Write a 50-word story about the individual inside the image." and the following system prompt, “You are a helpful chat assistant. You are going to generate texts in response to images depicting fictional individuals." [^3] The system prompt was not supplied to Llama-3.2 as it did not support system-level instructions. The maximum number of generated tokens was set to 150, and 50 stories were generated for each image, totaling 1,000 completions from each VLM. All data collection involving open-source models was conducted using an NVIDIA RTX A6000 GPU.
@@ 2.3 Measure of Homogeneity
To quantify homogeneity of stories generated for each group, we adopted the measure introduced by Lee et al. (2024). We first represented the generated stories into sentence embedding representations using a pre-trained Sentence-BERT model (Reimers and Gurevych, 2019)—specifically all-mpnet-base-v2—and then calculated the cosine similarity between all possible combinations of sentence embeddings of stories generated for each image. Larger cosine similarity between sentence embeddings indicates that the stories are more similar to each other and, hence, more homogeneous.
@@ 2.4 Comparison of Cosine Similarity Measures
We fitted mixed-effects models (Bates et al., 2014; Pinheiro and Bates, 2000) to compare cosine similarity measurements between groups. These models account for random variations in measurements, such as those arising from repeated cosine similarity values for texts generated from images within the same condition (i.e., skin tone and gender), as well as variations in baseline similarity across conditions. To account for the resemblance between facial stimuli generated from the same set thereby possibly affecting the similarity of the generated stories, the sets of facial stimuli (i.e. Pair ID) were used as random intercepts in all our mixed-effects models.
First, we fitted a Skin Tone Model with skin tone as the sole fixed effect to test the hypothesis that stories about darker-skinned Black individuals are more homogeneous than those about lighter-skinned Black individuals. Next, we fitted a Gender Model with gender as the sole fixed effect to test the hypothesis that stories about women are more homogeneous than those about men. Finally, we fitted a Skin Tone and Gender Interaction Model with skin tone, gender, and their interactions to examine the interaction between skin tone and gender.
In all models, lighter skin tone was set as the reference level, with a significantly positive effect of skin tone indicating larger cosine similarity values for darker-skinned Black individuals than for lighter-skinned Black individuals. For gender, men were set as the reference level, with a significantly positive effect indicating higher cosine similarity values for Black women compared to Black men. Mixed-effects models were fitted using the lme4 package (Lenth et al., 2024), and skin tone effects within gender groups were evaluated using the emmeans package (Lenth et al., 2024). Likelihood-ratio tests, conducted with the afex package, assessed whether adding individual terms improved model fit. All analyses were performed in R version 4.4.0.
2.5 Power Analysis
We used 10 facial stimuli per group to assess homogeneity bias. Power analysis with the simr package in R (Green et al., 2023), employing Monte Carlo simulations, determined that 1,245 cosine similarity measurements per pair of stimuli (i.e., Pair ID) were required to achieve 90% power for detecting an interaction effect (skin tone $\times$ gender) with an effect size of 0.30 (from Lee et al. (2024)) at $\alpha$ = .05. With 2,500 cosine similarity measurements collected, our study exceeded this requirement. This analysis accounted for the increased power requirements of detecting interaction effects, ensuring that our study was adequately powered to test all effects with statistical confidence.
@@ 3 Results
In the Results section, we summarize the Skin Tone Model output to evaluate the effect of skin tone, presenting likelihood-ratio test results and a visualization of cosine similarity measurements for each skin tone group. We then summarize the Gender Model output to evaluate the effect of gender, presenting likelihood-ratio test results and a visualization of cosine similarity measurements for each gender group. Finally, we analyze the Interaction term from the Skin Tone and Gender Interaction Model, presenting likelihood-ratio test results, simple slopes, and a visualization of cosine similarity measurements for each intersectional group.
3.1 Darker Skin Tone Associated with Increased Homogeneity
Figure 3: Standardized cosine similarity values of lighter and darker skinned Black individuals generated from all four VLMs. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
As shown in Figure 3, stories about darker-skinned Black individuals were significantly more homogeneous than those about lighter-skinned Black individuals in GPT-4o mini, GPT-4 Turbo, and Llama-3.2 (bs = 0.044, 0.15, and 0.080, respectively, ps < .001). However, we found an opposite pattern in BLIP-3 (bs = -0.095, ps < .001). Likelihood-ratio tests revealed that including skin tone improved model fit for all four VLMs ($\chi^{2}$(1)s $\geq$ 289.55, ps < .001). See Table S2 for summary output of the Skin Tone Models and Table S5 for likelihood-ratio test results.

### v2

# Vision-Language Models Generate More Homogeneous Stories for Phenotypically Black Individuals

@@ Abstract
Vision-Language Models (VLMs) extend Large Language Models’ capabilities by integrating image processing, but concerns persist about their potential to reproduce and amplify human biases. While research has documented how these models perpetuate stereotypes across demographic groups, most work has focused on between-group biases rather than within-group differences. This study investigates homogeneity bias—the tendency to portray groups as more uniform than they are—within Black Americans, examining how perceived racial phenotypicality influences VLMs’ outputs. Using computer-generated images that systematically vary in phenotypicality, we prompted VLMs to generate stories about these individuals and measured text similarity to assess content homogeneity. Our findings reveal three key patterns: First, VLMs generate significantly more homogeneous stories about Black individuals with higher phenotypicality compared to those with lower phenotypicality. Second, stories about Black women consistently display greater homogeneity than those about Black men across all models tested. Third, in two of three VLMs, this homogeneity bias is primarily driven by a pronounced interaction where phenotypicality strongly influences content variation for Black women but has minimal impact for Black men. These results demonstrate how intersectionality shapes AI-generated representations and highlight the persistence of stereotyping that mirror documented biases in human perception, where increased racial phenotypicality leads to greater stereotyping and less individualized representation.
@@ 1 Introduction
1.1 Homogeneity bias in artificial intelligence
In addition to those biases, AI systems also demonstrate more subtle forms of stereotyping, specifically homogeneity bias—a tendency to portray certain groups with less individuality and variation than others. This phenomenon relates to the perceived variability literature in social psychology, which examines how certain groups are represented as more similar to one another than others (Linville et al., 1989; Quattrone and Jones, 1980).
Recent research has documented this phenomenon in language models. Lee et al. (2024) found that ChatGPT generated more uniform texts for racial/ethnic minorities and women compared to White Americans and men, suggesting this stems from imbalanced representation and stereotypical portrayals in training data. In parallel work, Cheng et al. (2023) showed how AI-generated texts about marginalized groups often amplify defining characteristics, creating caricature-like representations rather than nuanced portrayals of individuals. These findings highlight a concerning pattern in how AI systems process and reproduce information about diverse social groups.
@@ 1.1 Homogeneity bias in artificial intelligence
1.2 The effect of racial phenotypicality on stereotyping
Most work examining bias in AI systems focus on between-group biases (e.g., whether Black people are more associated with negative traits than White people), while neglecting within-group differences. Research in social psychology, however, has extensively documented that individuals within the same racial group can experience different degrees of bias based on their physical characteristics. Racial phenotypicality refers to the degree to which a person’s physical features are perceived as typical of their racial group. For Black individuals, these features include skin tone, hair texture, lip thickness, and nose width, among others (Hagiwara et al., 2012; Stepanova and Strube, 2012). Studies show that Black individuals who are perceived as having more typically Black features experience greater stereotyping than those with less typical features (e.g., Stepanova and Strube, 2018; Kahn and Davies, 2011; Maddox, 2004). This is often referred to as racial phenotypicality bias.
This bias manifests in significant real-world consequences: Black individuals with higher perceived racial phenotypicality receive lower ratings and fewer job offers in hiring scenarios (Wade et al., 2004; Harrison and Thomas, 2009), achieve lower levels of educational attainment and income (Keith and Herring, 1991), experience greater racial discrimination (Klonoff and Landrine, 2000), and report higher levels of mental distress due to discrimination (Gleiberman et al., 1995). These findings underscore how phenotypicality plays a critical role in shaping both perceptions and life outcomes for Black individuals in the United States.
Despite substantial research on how perceived racial phenotypicality affects social perceptions and outcomes, few studies have explored this phenomenon in Artificial Intelligence (AI) models, particularly in Vision-Language Models (VLMs). Earlier work by Buolamwini and Gebru (2018) revealed that commercial gender-classification systems perform significantly better for lighter-skinned individuals, with more pronounced disparities for women than men. While this research focused specifically on skin tone–only one component of racial phenotypicality rather than the full range of phenotypic features–it highlighted disparities in AI performance based on physical characteristics. Subsequent research has investigated the effect of skin tone on machine learning model performance (e.g., Groh et al., 2024; Kinyanjui et al., 2019), but the relationship between the full spectrum of racial phenotypic features and bias in newer generative models remains relatively under-explored, especially regarding how it might affect the homogeneity of content generated about individuals.
@@ 1.3 This work
We investigate how phenotypicality influences homogeneity bias in VLMs by examining whether higher levels of phenotypicality are associated with greater uniformity in generated content. Drawing from social psychology literature on racial phenotypicality effects, we hypothesized that VLMs would produce more homogeneous stories about individuals with higher racial phenotypicality. This approach moves beyond traditional between-group comparisons to examine within-group effects, addressing a critical gap in AI bias research.
Figure 1: Summary of the experimental setup. We collect 50-word stories about Black individuals differing in phenotypicality (i.e., lower and higher phenotypicality) using four state-of-the-art Vision-Language Models. After encoding these stories into sentence embeddings, we compare the pairwise similarity of the embeddings using mixed-effects models.
This work builds upon concurrent work Lee et al. (2025) that found no significant relationship between racial phenotypicality and homogeneity bias in VLMs, though they observed that gender phenotypicality was associated with increased homogeneity. Our approach differs methodologically through the use of computer-generated stimuli that enable more controlled manipulation of phenotypicality without confounding variables present in the real-world images they used.
@@ 2 Method
We first explain our process for selecting images representing Black men and women with lower and higher racial phenotypicality. Next, we detail our VLM selection criteria and the prompts used for data collection. Finally, we describe our methodology for measuring and comparing pairwise similarity between the stories generated for these images.
@@ 2.1 Image stimuli
We sampled ten image sets of Black American men and women from the publicly available GAN Face Database (GANFD; Marsden et al., 2024), which features realistic, computer-generated faces. The database includes sets of images representing the same fictional individuals, with manipulations applied to vary facial features associated with perceived race, specifically those that influence racial phenotypicality.
Our selection process for stimuli only considered images where the face was categorized as either "Black" or "Multiple" race based on human ratings.111”Black” categorization indicates that over 50% of human raters categorized the face as Black/African American, while ”Multiple” indicates either no category reached the 50% threshold or the top two categories were within 10 percentage points of each other. For each set, we applied one of three selection strategies based on the available images. If a set contained more than two images categorized as "Black," we selected the images with the highest and lowest perceived Blackness ratings (measured on a 0-100 scale). If a set contained one "Black" image and at least one "Multiple" race image, we selected the "Black" image and the "Multiple" race image with the highest perceived Blackness rating. If a set contained only "Multiple" race images, we selected the two images with the highest perceived Blackness ratings. Only sets yielding exactly two images were included in our final stimulus selection. Within each resulting pair, we labeled the image with the lower perceived Blackness score as "lower phenotypicality" and the image with the higher score as "higher phenotypicality." This methodological approach provided experimental control while ensuring meaningful phenotypicality differences within each pair. Finally, to ensure consistency, we used cropped images with a uniform grey background that contained only the face, allowing us to isolate the effect of phenotypicality while holding other visual characteristics constant. See Figure 2 for sample image pairs used in the study.
Figure 2: Two face stimulus pairs representing (A) Black men and (B) Black women. In each pair, the left image depicts a Black individual with lower phenotypicality, while the right shows a Black individual with higher phenotypicality, generated from the same set.
@@ 2.2 Selection of vision-language models and writing prompts
We used a set of Vision-Language Models capable of processing facial images to write stories. [^2] See Section S2 for details on why certain models such as BLIP-3 (Xue et al., 2024) and Claude 3.7 Sonnet (Anthropic, 2025) were excluded from our study. Our analysis included two proprietary VLMs—GPT-4o mini and GPT-4 Turbo—and an open-source VLM—Llama-3.2 (Llama-3.2-11B-Vision-Instruct; Grattafiori et al., 2024). We accessed the proprietary models using the OpenAI API and the open-source models by downloading the model weights and running inferences on them locally.
The models were given the following writing prompt, “Write a 50-word story about the individual inside the image." and the following system prompt, “You are a helpful chat assistant. You are going to generate texts in response to images depicting fictional individuals." [^3] The system prompt was not supplied to Llama-3.2 as it did not support system-level instructions. The maximum number of generated tokens was set to 150. We used 10 facial stimuli per group to assess homogeneity bias. Based on power analysis with the simr package in R (Green et al., 2023), we determined that 1,245 cosine similarity measurements per pair of stimuli (i.e., Pair ID) were required to achieve 90% power for detecting an interaction effect (phenotypicality $\times$ gender) with an effect size of 0.30 (from Lee et al. (2024)) at $\alpha$ = .05. To satisfy this requirement, we collected 50 stories for each of the images, totaling 2,500 cosine similarity measurements per Pair ID. This approach ensured that our study was adequately powered to test all effects with statistical confidence. All data collection involving open-source models was conducted using an NVIDIA RTX A6000 GPU.
@@ 2.3 Measure of homogeneity
To quantify homogeneity of stories generated for each group, we adopted the measure introduced by Lee et al. (2024). We first represented the generated stories into sentence embedding representations using a pre-trained Sentence-BERT model (Reimers and Gurevych, 2019)—specifically all-mpnet-base-v2—and then calculated the cosine similarity between all possible combinations of sentence embeddings of stories generated for each image. Larger cosine similarity between sentence embeddings indicates that the stories are more similar to each other and, hence, more homogeneous. This embedding-based approach for measuring text similarity has become standard practice in social science research (Lin, 2025; Licht, 2023), as it effectively captures semantic relationships between texts even when they use different vocabulary to express similar meanings, a significant advantage over traditional word-overlap methods. We present examples of text pairs with varying degrees of similarity in Table 1, which demonstrates the face validity of our measure as texts become noticeably less similar in content and narrative as the cosine similarity value decreases.
Table 1: Examples of text pairs generated by GPT-4o mini for Black women with higher phenotypicality, ordered by percentile of cosine similarity. This progression provides face validity for the cosine similarity measure, as texts become noticeably less similar in content and narrative as the cosine similarity value decreases.
Percentile Text 1 Text 2 Cos. Sim.
0th In a bustling city, Maya started a community garden, bringing neighbors together. Each weekend, laughter echoed as they planted seeds and shared recipes. With every sprout, friendships blossomed. Maya’s vision transformed a barren lot into a vibrant oasis… In a bustling city, Maya launched a community garden, uniting neighbors through shared soil and laughter. Each seed planted blossomed with hope, mirroring her radiant smile. As fruits ripened, so did friendships… 0.97
25th In a quiet village, Maya unearthed forgotten stories. With every word, she brought history alive, reuniting families with their lost legacies. One day, she stumbled upon her own family’s past, discovering a treasure of love letters… Under the bustling city lights, Maya dreamed of painting the world in vibrant colors. Each stroke on her canvas reflected her journey—filled with hope, struggles, and laughter. One evening, inspired by the sunset… 0.65
50th In a bustling city, Maya found joy in small moments—sharing laughs with friends, volunteering at the local shelter, and painting sunsets that mirrored her dreams. With her infectious smile, she lit up every room… In a bustling city, Mira cherished simple moments—coffee with her grandmother, sunsets by the river. One day, she found a forgotten book in a tiny shop, its pages filled with stories of adventure… 0.55
75th In a small town, Malik found joy in simple moments: the sunrise over the hills, children laughing at the park, and the aroma of fresh coffee in the morning. A skilled painter, he captured these scenes on canvas… In a bustling city, Alex discovered an old guitar in his grandfather’s attic. He spent nights learning chords, transforming his emotions into melodies. One rainy evening, he played at a local café… 0.46
100th In a small town, Marcus discovered an ancient map while renovating his grandmother’s attic. Intrigued, he embarked on a weekend adventure. The map led him to a hidden waterfall, where he found a forgotten journal… In a quiet café, Raj scribbled ideas for his next invention. His passion for technology sparked a dream: a device to help others communicate effortlessly. With each stroke of his pen, he envisioned a world… 0.12
@@ 2.4 Comparison of cosine similarity measures
We fitted three mixed-effects models (Bates et al., 2014; Pinheiro and Bates, 2000) to compare cosine similarities across groups for each VLM. These models account for random variations in measurements while controlling for image pair effects. To account for the resemblance between facial stimuli generated from the same set thereby possibly affecting the similarity of the generated stories, the sets of facial stimuli (i.e. Pair ID) were used as random intercepts in all our mixed-effects models.
First, we fitted a Phenotypicality model with phenotypicality as the sole fixed effect to test the hypothesis that stories about Black individuals with higher phenotypicality are more homogeneous than those about Black individuals with lower phenotypicality. In this model, lower phenotypicality was set as the reference level, with a significantly positive effect of phenotypicality indicating larger cosine similarity values for Black individuals with higher phenotypicality.
Next, we fitted a Gender Model with gender as the sole fixed effect to test the hypothesis that stories about women are more homogeneous than those about men. In this model, men were set as the reference level, with a significantly positive gender effect indicating higher cosine similarity values for Black women compared to Black men.
Finally, we fitted an Interaction Model with phenotypicality, gender, and their interactions to examine the interaction between phenotypicality and gender. In this model, the phenotypicality term represents the effect of phenotypicality for men (the reference gender group), the gender term represents the effect of gender for Black individuals with lower perceived racial phenotypicality (the reference phenotypicality group), and the interaction term indicates how the effect of phenotypicality differs for women compared to men.
The models were fitted using the lme4 package (Bates et al., 2024), and phenotypicality effects within gender groups were evaluated using the emmeans package (Lenth et al., 2024). Likelihood-ratio tests, conducted with the afex package (Singmann et al., 2024), assessed whether adding individual terms improved model fit. All analyses were performed in R version 4.4.0.
@@ 3 Results
In the Results section, we summarize the Phenotypicality Model output to evaluate the effect of phenotypicality, presenting likelihood-ratio test results and a visualization of cosine similarity measurements for each phenotypicality group. We then summarize the Gender Model output to evaluate the effect of gender, presenting likelihood-ratio test results and a visualization of cosine similarity measurements for each gender group. Finally, we analyze the Interaction term from the Interaction Model. We present likelihood-ratio test results comparing models with and without the interaction. We then conduct simple slopes analysis to examine the effect of phenotypicality within each gender group. To visualize these findings, we present plots of cosine similarity measurements for each intersectional group.
3.1 Higher phenotypicality associated with increased homogeneity
Figure 3: Standardized cosine similarity values of Black individuals with lower versus higher phenotypicality ratings generated from all four VLMs. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
Stories about Black individuals with higher phenotypicality were significantly more homogeneous than those about Black individuals with lower phenotypicality in all VLMs (bs = 0.044, 0.15, and 0.080, respectively, ps < .001; see Figure 3). Likelihood-ratio tests revealed that including phenotypicality improved model fit for all VLMs ($\chi^{2}$(1)s $\geq$ 289.55, ps < .001). See Table S2 for summary output of the Phenotypicality Models and Table S5 for likelihood-ratio test results.